# Early Directional Convergence in Deep Homogeneous Neural Networks for Small Initializations

### v1

@@ 5.1.2 A Closer Look at the Weights
In Theorem 5.1 we established that in the early stages of training, for sufficiently small initialization, the weights of neural network approximately converge in direction along the KKT points of the NCF while their norm remained small. In this section, we use a toy example to highlight the emergence of additional structural properties in the weights during the early stage of training that are not explained by Theorem 5.1. We train the following $4-$homogeneous neural network $\mathcal{H}(\mathbf{x};\mathbf{W}_{1},\mathbf{W}_{2})=\mathbf{1}^{\top}\max(\mathbf{W}_{2}\mathbf{W}_{1}\mathbf{x},\mathbf{0})^{2}$, where $\mathbf{W}_{1}\in\mathbb{R}^{20\times 10}$, $\mathbf{W}_{2}\in\mathbb{R}^{20\times 20}$ and $\mathbf{1}\in\mathbb{R}^{20}$ is the all-one vector. Here, we train to minimize the square loss with respect to the output of a smaller neural network *$\mathcal{H}^{*}(\mathbf{x};\mathbf{W}_{1}^{*},\mathbf{W}_{2}^{*})=\mathbf{1}^{\top}\max(\mathbf{W}_{2}^{*}\mathbf{W}_{1}^{*}\mathbf{x},\mathbf{0})^{2}$, where the entries of $\mathbf{W}_{1}^{*}\in\mathbb{R}^{2\times 10}$, $\mathbf{W}_{2}^{*}\in\mathbb{R}^{3\times 2}$ are drawn from the standard normal distribution and $\mathbf{1}\in\mathbb{R}^{3}$ is the all-one vector. The training data has 100 points sampled uniformly from unit sphere in $\mathbb{R}^{10}$. Let $\mathbf{w}$ be a vector containing all the entries of $\mathbf{W}_{1},\mathbf{W}_{2}$, and $\tilde{\mathbf{w}}=\mathbf{w}/\|\mathbf{w}\|_{2}.$ For training, we use gradient descent with step-size $5\cdot 10^{-3}$, and the initialization $\mathbf{w}(0)=\delta\mathbf{w}_{0}$, where $\delta=0.1$ and $\mathbf{w}_{0}$ is a random unit norm vector. We train for 99100 iterations, and in Figure 0(a), plot the evolution of training loss and $\ell_{2}-$norm of the weights. It is evident from the figure that loss almost remains the same, and the norm of the weights has increased but only by a constant factor and is still small. It can be easily shown (see Lemma B.3) that a unit norm vector is a KKT point of the constrained NCF if the vector and the gradient of the NCF at that vector are parallel. Therefore, to measure the directional convergence of $\mathbf{w}$ to the KKT point of the constrained NCF, in Figure 0(b) we plot the evolution of the inner product between $\tilde{\mathbf{w}}(t)$ and $\nabla\mathcal{N}(\tilde{\mathbf{w}}(t))/\|\nabla\mathcal{N}(\tilde{\mathbf{w}}(t))\|_{2}$, where $\mathcal{N}(\cdot)$ denotes the NCF for this example. Clearly, in accordance with Theorem 5.1, the weights have approximately converged in direction to a KKT point of the constrained NCF.
@@
Figure 1: Early training dynamics of a squared-ReLU neural network. Panel $(a)$: the evolution of training loss (normalized with respect to loss at initialization) and the $\ell_{2}$-norm of all the weights with iterations. Panel $(b)$: the evolution of $\nabla\mathcal{N}(\tilde{\mathbf{w}}(t))^{\top}\tilde{\mathbf{w}}(t)/\|\nabla\mathcal{N}(\tilde{\mathbf{w}}(t))\|_{2}$ (a measure of directional convergence of $\mathbf{w}(t)$). We observe that the loss has barely changed, and the norm of weights has increased by a constant factor but remains small. Also, the weights approximately converge in direction to a KKT point of the constrained NCF.
We next take a closer look at the weights of the neural networks at iteration 99100, i.e., when we stop training and the weights have approximately converged in direction to a KKT point of the constrained NCF. In Figure 2, we plot each row of $\mathbf{W}_{1}/\|\mathbf{w}\|_{2}$ and $\mathbf{W}_{2}/\|\mathbf{w}\|_{2}$ at initialization and at iteration 99100. At initialization the rows of $\mathbf{W}_{1}/\|\mathbf{w}\|_{2}$ and $\mathbf{W}_{2}/\|\mathbf{w}\|_{2}$ appear to be randomly oriented, as expected. At iteration 99100, all the rows seem to be oriented along certain specific directions and there is a low-rank structure among $\mathbf{W}_{1}/\|\mathbf{w}\|_{2}$ and $\mathbf{W}_{2}/\|\mathbf{w}\|_{2}$. The low-rank structure can be further verified by observing the singular values of $\mathbf{W}_{1}/\|\mathbf{w}\|_{2}$ and $\mathbf{W}_{2}/\|\mathbf{w}\|_{2}$, depicted in Figure 3.
Figure 2: Each curve in the plots correspond to a row of $\mathbf{W}_{1}/\|\mathbf{w}\|_{2}$ or $\mathbf{W}_{2}/\|\mathbf{w}\|_{2}$ at different stages of training. Panel $(a)$: $\mathbf{W}_{1}/\|\mathbf{w}\|_{2}$ at initialization. Panel $(b)$: $\mathbf{W}_{1}/\|\mathbf{w}\|_{2}$ at iteration 99100. Panel $(c)$: $\mathbf{W}_{2}/\|\mathbf{w}\|_{2}$ at initialization. Panel $(d)$: $\mathbf{W}_{2}/\|\mathbf{w}\|_{2}$ at iteration 99100.
Figure 3: Panel $(a)$: singular values of $\mathbf{W}_{1}/\|\mathbf{w}\|_{2}$ at iteration 99100. Panel $(b)$: singular values of $\mathbf{W}_{2}/\|\mathbf{w}\|_{2}$ at iteration 99100.
We clarify that the results of Theorem 5.1 neither explain how such low-rank structures emerge in the weights, nor characterizes such low-rank structures. We leave it as an important future research direction to explain the emergence of these low-rank structures. Notably, similar low-rank behavior under small initialization was observed by [29] for three-layer ReLU networks (which is 3-homogeneous). Also, [30] has empirically shown emergence of low-rank structures in the kernel learned by neural networks during the early phase of training.
5.1.3 Challenges for Non-differentiable Neural Networks

### v2

@@ 4.1.2 A Closer Look at the Weights
Figure 1: Early training dynamics of a squared-ReLU neural network. Panel $(a)$: the evolution of training loss (normalized with respect to loss at initialization) and the $\ell_{2}$-norm of all the weights with iterations. Panel $(b)$: the evolution of $\nabla\mathcal{N}(\widetilde{\mathbf{w}}(t))^{\top}\widetilde{\mathbf{w}}(t)/\|\nabla\mathcal{N}(\widetilde{\mathbf{w}}(t))\|_{2}$ (a measure of directional convergence of $\mathbf{w}(t)$). Note that the loss has barely changed, and the norm of weights has increased by a constant factor but remains small. Also, the weights approximately converge in direction to a KKT point of the constrained NCF.
In this section, we use a toy example to highlight the emergence of additional structural properties in the weights during the early stage of training that are not explained by Theorem 1.